### v1

@@ 5.1.1 Finite-energy functions and Dirichlet truncation
We now claim that the limits $h_{1},\ldots,h_{k}$ are linearly independent. Suppose for contradiction that $\dim\left(\operatorname{Span}\left\{h_{1},\ldots,h_{k}\right\}\right)\leq k-1$. Then there exists an eigenfunction $\widehat{h}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ associated with $\lambda_{\ell}\left(A^{\mathrm{D}}\right)$, $\ell\leq k$ that is linearly independent of $h_{1},\ldots,h_{k}$. As in the proof of Lemma 5.3, we approximate $\widehat{h}$ by a smooth function $\mathcal{U}$ with $\operatorname{supp}(\mathcal{U})\subset B_{R_{m}}$ and consider
$M:=\operatorname{Span}\left\{u_{1,R_{m}},\ldots,u_{k,R_{m}},\mathcal{U}\right\}.$
For sufficiently large $m$, we have $\dim(M)=k+1$. Hence, using the variational characterisation,
$\sigma_{k+1}^{\mathrm{D}}(\Omega^{\mathrm{ext}}_{R_{m}})\leq\sup_{u\in M,\,u\neq 0}\frac{\int_{\Omega^{\mathrm{ext}}_{R_{m}}}|\nabla u|^{2}\,\mathrm{d}x}{\int_{\partial\Omega}|u|^{2}\,\mathrm{d}S}.$
As in the proof of Lemma 5.3, the right-hand side converges to $\lambda_{k}\left(A^{\mathrm{D}}\right)$ as $m\to\infty$. If we consider a $k$ where $\lambda_{k}\left(A^{\mathrm{D}}\right)\neq\lambda_{k+1}\left(A^{\mathrm{D}}\right)$, this forms a contradiction. Consequently, for any eigenfunction $f_{q}$ associated with $\lambda_{q}\left(A^{\mathrm{D}}\right)$, we can set $k=\max\left\{p\in\mathbb{N}:\lambda_{p}\left(A^{\mathrm{D}}\right)\leq\lambda_{q}\left(A^{\mathrm{D}}\right)\right\}$ and obtain the existence of an $r\leq k$ with $h_{r}=f_{q}$. By definition, the sequence $u_{r,R_{m}}$} satisfies the required properties. ∎
@@ 5.1.2 Finite-energy functions and Helmholtz equation
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and let $k\in\mathbb{N}$. Then,
@@
Define the $k$-dimensional subspace $M:=\operatorname{Span}\left\{g_{1},\ldots,g_{k}\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$. Using the variational characterisation of $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$, we obtain
@@
We now proceed by induction on $k$. Let $m(k)$ denote the multiplicity of $\lambda_{k}\left(A^{\mathrm{D}}\right)$, and define $M(k):=\sum_{j=1}^{k}m(k)$. Then
@@ 5.1.3 Vanishing flow and Neumann truncation
In §4.3.2 we introduced, via layer potentials, an alternative formulation of the exterior Steklov problem. In this setting, functions are not required to vanish at infinity, but their gradients must decay faster than $|x|^{1-n}$. We now prove that this formulation is equivalent to the Neumann truncation.
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and let $k\in\mathbb{N}$. Denote by $\chi_{k}\left(\Omega^{\mathrm{ext}}\right)$ the eigenvalues of ($\mathrm{VF}$), and recall that $\sigma^{\mathrm{N}}_{k}\left(\Omega^{\mathrm{ext}}_{R}\right)$ are the eigenvalues of ($\mathrm{SN}_{R}$) and. Then
$\lim_{R\to\infty}\sigma^{\mathrm{N}}_{k}\left(\Omega^{\mathrm{ext}}_{R}\right)=\chi_{k}\left(\Omega^{\mathrm{ext}}\right).$
@@
By Theorem 3.9 (see also [AretE15, Theorem 5.6]), the resolvent of the Dirichlet-to-Neumann operator with Neumann condition on $\partial B_{R}$ converges uniformly to the resolvent of $A^{\mathrm{N}}$ from Definition 3.1. Thus it suffices to show the equivalence between $A^{\mathrm{N}}$ and ($\mathrm{VF}$).
Let $\left(\sigma^{\mathrm{N}},u^{\mathrm{N}}\right)\in\mathbb{R}\times W\left(\Omega^{\mathrm{ext}}\right)$ be an eigenpair of $A^{\mathrm{N}}$. Let $\langle u\rangle$ denote the average of $u$, as defined in (3.1). Then, by definition, $w:=u^{\mathrm{N}}-\left\langle u^{\mathrm{N}}\right\rangle$ is in $W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$, and therefore, by (3.4), $w\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. Without loss of generality, assume $\Omega\subset B_{1}$. As in the proof of Proposition 6.2, which will be presented later, we expand $w$ in spherical harmonics
@@
where ${u}_{(k,i)}\in E^{1}\left(B_{1}^{\mathrm{ext}}\right)$, given by ${u}_{(k,i)}(r,\theta)=r^{2-n-k}Y_{k,i}(\theta)$, are Steklov eigenfunctions in the exterior of the unit ball forming an orthonormal basis of $L^{2}\left(\partial B_{1}\right)$, see §2.2. This expansion converges uniformly in $\Omega^{\mathrm{ext}}$, see the proof of Proposition 6.2. From the explicit form of ${u}_{(k,i)}$ it follows that $w$ (and thus $u^{\mathrm{N}}$) satisfies the vanishing flow condition (4.9) if and only if $\left(w,{u}_{(0,1)}\right)_{L^{2}\left(\partial B_{1}\right)}=0$.
To establish this, note first that since $w$ is harmonic and $\partial_{\nu}w=\partial_{\nu}u^{\mathrm{N}}=\sigma^{\mathrm{N}}u^{\mathrm{N}}$ on $\partial\Omega$, we have
$0=\int_{\Omega^{\mathrm{ext}}\cap B_{1}}\Delta w\,\mathrm{d}x=\int_{\partial\Omega}\partial_{\nu}w\,\mathrm{d}S+\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\sigma^{\mathrm{N}}\int_{\partial\Omega}u^{\mathrm{N}}\,\mathrm{d}S+\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S.$
In the last step we used that the first eigenvalue vanishes and the associated eigenfunction is a constant and therefore, any other eigenfunction is orthogonal to a constant. Furthermore, $\partial_{\nu}$ coincides with the radial derivative on $\partial B_{1}$. Hence,
@@
Conversely, if $(\chi,u)$ is an eigenpair of ($\mathrm{VF}$), then $u\in W\left(\Omega^{\mathrm{ext}}\right)$ by definition, hence it is also an eigenpair of $A^{\mathrm{N}}$. ∎
@@ 5.2.1 Conformal mapping and truncated domains
By [AnPo21, Lemma 4.5], there exists a constant $C_{1}>0$ which depends only on the dimension such that any $f_{k,R}$ admits a harmonic extension $\widetilde{f}_{k,R}\in H^{1}(\Omega^{*})$ satisfying
@@
In analogy with Corollary 5.5, the convergence of the eigenvalues also implies convergence of the associated eigenfunctions in the following sense.
@@
We distinguish two cases. For the first part, we assume $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)\neq 0$. Let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ be an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})$, extended by zero to $\Omega^{\mathrm{ext}}$ and define
@@
If $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=0$, any eigenfunction is constant. Let $u_{k,R}$ be an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})$ normalised to $\|u_{k,R}\|_{L^{2}(\partial\Omega)}=1$. The convergence of the eigenvalue implies
@@
By the Poincaré inequality, this implies that $\left(u_{k,R}\circ\phi\right)-\frac{\int_{\Omega^{*}}u_{k,R}\circ\phi\,\mathrm{d}x}{|\Omega^{*}|}$ converges strongly to zero in $H^{1}(\Omega^{*})$. Moreover, since $\int_{\partial\Omega^{*}}|u_{k,R}\circ\phi|^{2}\,\mathrm{d}S\leq\frac{1}{\inf_{x\in\partial\Omega^{*}}|\phi^{\prime}(x)|}<\infty$ the average value of $u_{k,R}\circ\phi$ remains bounded. Therefore, there exists a subsequence that converges strongly in $H^{1}(\Omega^{*})$ to a constant $c\in\mathbb{R}$. ∎
@@
The case $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)=0$ can be handled as in Corollary 5.11. For $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)\neq 0$, we follow the proof of Theorem 5.10. Let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated to $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ normalised to $\|\nabla u_{k,R}\|_{L^{2}\left(\Omega^{\mathrm{ext}}_{R}\right)}=1$. Define
@@ 5.2.2 Conformal mapping and Helmholtz equation
In this subsection, we establish the equivalence between the formulation with conformal mappings and the formulation introduced in §3.3, based on the Helmholtz equation. Theorem  5.13 mirrors the convergence result in Theorem 5.9 for the Dirichlet problem.
@@
For the reverse inequality, let $f_{k}\in H^{1}(\Omega^{*})$ be an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$, normalised so that $\int_{\partial\Omega^{*}}|\phi|f_{j}f_{k}\,\mathrm{d}S=\delta_{j,k}$. We construct a cut-off function by
@@
To obtain a reverse inequality, define the $k$-dimensional spaces
@@ 5.2.3 Conformal mapping and layer potentials
Let $u_{k}\in H^{\frac{1}{2}}(\partial\Omega)$ be an eigenfunction associated with $\tau_{k}(\Omega^{\mathrm{ext}})$, and let $U_{k}$ be its harmonic extension. Then, $U_{k}\circ\phi$ coincides, up to a constant factor, with an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$.
@@
$\left|f_{k}(x)-w_{k}(0)\right|=\left|w_{k}\left(\frac{x}{|x|^{2}}\right)-w_{k}(0)\right|=\frac{|\nabla w_{k}(x_{0})|}{|x|^{2}}$
@@
is a solution of ($\mathrm{CT}$). That completes the proof. ∎
@@
Let $n\geq 3$, let $u$ be a solution of ($\mathrm{ES}$) and let $f$ be the restriction of $u$ to $\partial\Omega$. It follows from [AuHa14b, Section 5] that there is a unique $U\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ such that $U=f$ on $\partial\Omega$. Moreover, $U=O(|x|^{2-n})$ by [Xio23, Proposition 12] (see also Proposition 6.1). It follows that UuU-u is harmonic, vanishing on $\Gamma$ and decaying at $\infty$, whence $u=U\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. So, any solution of ($\mathrm{ES}$) is in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$. It then follows by integration by parts against test function that $u$ is a solution of the finite energy weak Steklov eigenvalue equation (3.5). On the other hand, Theorem 3.5 provides a sequence of weak Steklov eigenfunction $u_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ whose restriction to $\partial\Omega$ form an orthogonal basis of $L^{2}(\partial\Omega)$. It follows from [Xio23, Proposition 12] that these functions satisfy (1.2), and hence they are solutions of ($\mathrm{ES}$).
Let $n=2$ and let $u$ be a solution of ($\mathrm{ES}$). Because $u$ is harmonic and bounded, it converges to a finite value at infinity, denoted by $u_{\infty}$. With $\phi$ as is (3.10), we consider the function $v:\Omega^{*}\to\mathbb{R}$, $v=u\circ\phi$ in $\Omega^{*}\setminus\{0\}$ and $v(0)=u_{\infty}$. Then, $v$ is harmonic and satisfies ($\mathrm{CT}$) and so it is a Steklov eigenfunction in the conformal sense. Similarly, any solution of ($\mathrm{CT}$) will give a solution of ($\mathrm{ES}$) once it is inverted through $\phi$. ∎

### v2

@@ 5.1.1 Finite-energy functions and Dirichlet truncation
Since
$\int_{\partial\Omega}h_{i}h_{j}\,\mathrm{d}S=\lim_{m\to\infty}\int_{\partial\Omega}u_{i,R_{m}}u_{j,R_{m}}\,\mathrm{d}S=\lim_{m\to\infty}\frac{\int_{\Omega^{\mathrm{ext}}_{R_{m}}}\left\langle\nabla{u}_{i,R_{m}},\nabla{u}_{j,R_{m}}\right\rangle\,\mathrm{d}x}{\sigma_{j}^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}_{R_{m}}\right)}=\delta_{i,j}\frac{1}{\lambda_{j}\left(A^{\mathrm{D}}\right)},$
the limits $h_{1},\ldots,h_{k}$ are linearly independent. Consequently, for any eigenfunction $f_{q}$ associated with $\lambda_{q}\left(A^{\mathrm{D}}\right)$, we can set $k=\max\left\{p\in\mathbb{N}:\lambda_{p}\left(A^{\mathrm{D}}\right)\leq\lambda_{q}\left(A^{\mathrm{D}}\right)\right\}$ and obtain the existence of an $r\leq k$ with $h_{r}=f_{q}$. By definition, the sequence $u_{r,R_{m}}$} satisfies the required properties. ∎
@@ 5.1.2 Finite-energy functions and Helmholtz equation
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$, and let $k\in\mathbb{N}$. Then,
@@
Define $M:=\operatorname{Span}\left\{g_{1},\ldots,g_{k}\right\}\subset H^{1}\left(\Omega^{\mathrm{ext}}\right)$. Since the restrictions of $g_{m}$ to $\partial\Omega$ are part of a basis, $\operatorname{dim}(M)=k$. Using the variational characterisation of $\mu_{k}\left(\Omega^{\mathrm{ext}},\Lambda\right)$, we obtain
@@
We now proceed by induction on $k$. Let
$M(k):=\left|\left\{j\in\mathbb{N}:\lambda_{j}\left(A^{\mathrm{D}}\right)\leq\lambda_{k}\left(A^{\mathrm{D}}\right)\right\}\right|.$
Then
@@ 5.1.3 Vanishing flow and Neumann truncation
In §4.3.2 we introduced, via layer potentials, an alternative formulation of the exterior Steklov problem. In this setting, functions are not required to vanish at infinity, but their gradients must decay faster than $|x|^{1-n}$. We now prove that this formulation is equivalent to the Neumann truncation. Recall that the eigenvalues $\lambda_{k}\left(A^{\mathrm{N}}\right)$ of the operator $A^{\mathrm{N}}$, defined in §3.2, satisfy
$\lambda_{k}\left(A^{\mathrm{N}}\right)=\lim_{R\to\infty}\sigma^{\mathrm{N}}_{k}\left(\Omega^{\mathrm{ext}}_{R}\right),\qquad k\in\mathbb{N}.$
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded open set with Lipschitz boundary and with connected $\Omega^{\mathrm{ext}}$. Then the spectrum of ($\mathrm{VF}$) is discrete, and eigenvalues of ($\mathrm{VF}$) coincide with those of $A^{\mathrm{N}}$ with account of multiplicities.
@@
Let $\lambda^{\mathrm{N}}\in\mathbb{R}$ be an eigenvalue of $A^{\mathrm{N}}$, and let $u^{\mathrm{N}}\in W\left(\Omega^{\mathrm{ext}}\right)$ be such that the trace of $u^{\mathrm{N}}$ is an eigenfunction associated with $\lambda^{\mathrm{N}}$, that is, $\Delta u^{\mathrm{N}}=0$ and $\partial_{\nu}u^{\mathrm{N}}=A^{\mathrm{N}}\left(\left.u^{\mathrm{N}}\right|_{\partial\Omega}\right)$. Let $\left\langle u^{\mathrm{N}}\right\rangle$ denote the average of $u^{\mathrm{N}}$, as defined in (3.1). Then, by definition, $w:=u^{\mathrm{N}}-\left\langle u^{\mathrm{N}}\right\rangle$ is in $W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right)$, and therefore, by (3.4), $w\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. Without loss of generality, assume $\Omega\subset B_{1}$. As in the proof of Proposition 6.2, which will be presented later, we expand $w$ in spherical harmonics
@@
where ${u}_{(k,i)}\in E^{1}\left(B_{1}^{\mathrm{ext}}\right)$, given by ${u}_{(k,i)}(r,\theta)=r^{2-n-k}Y_{k,i}(\theta)$, are Steklov eigenfunctions in the exterior of the unit ball whose traces form an orthonormal basis of $L^{2}\left(\partial B_{1}\right)$, see §2.2. Following the proof of Proposition 6.2 and using the decay at infinity, we obtain that this expansion converges uniformly in $\Omega^{\mathrm{ext}}$.
From the explicit form of ${u}_{(k,i)}$ it follows that $w$ (and thus $u^{\mathrm{N}}$) satisfies the vanishing flow condition (4.9) if and only if $\left(w,{u}_{(0,1)}\right)_{L^{2}\left(\partial B_{1}\right)}=0$. To establish this, note first that since $w$ is harmonic and $\partial_{\nu}w=\partial_{\nu}u^{\mathrm{N}}=\lambda^{\mathrm{N}}u^{\mathrm{N}}$ on $\partial\Omega$, we have
$0=\int_{\Omega^{\mathrm{ext}}\cap B_{1}}\Delta w\,\mathrm{d}x=\int_{\partial\Omega}\partial_{\nu}w\,\mathrm{d}S+\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\lambda^{\mathrm{N}}\int_{\partial\Omega}u^{\mathrm{N}}\,\mathrm{d}S+\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S=\int_{\partial B_{1}}\partial_{\nu}w\,\mathrm{d}S.$
In the last step we used that the first eigenvalue vanishes and the associated eigenfunction is a constant, and, therefore, any other eigenfunction is orthogonal to a constant. Furthermore, $\partial_{\nu}$ coincides with the radial derivative on $\partial B_{1}$. Hence,
@@
Conversely, if $(\chi,U)$ is an eigenpair of ($\mathrm{VF}$), then $U\in W\left(\Omega^{\mathrm{ext}}\right)$ by definition, hence $\chi$ and $U|_{\partial\Omega}$ are an eigenvalue and eigenfunction of $A^{\mathrm{N}}$. ∎
@@ 5.2.1 Conformal mapping and truncated domains
By [AnPo21, Lemma 4.5], there exists a constant $C_{1}>0$ which depends only on the dimension such that any $f_{k,R}$ admits an extension $\widetilde{f}_{k,R}\in H^{1}(\Omega^{*})$ which is harmonic in $B_{1/R}$, satisfying
@@
We proceed, by analogy with Corollary 5.5, to infer the convergence of the eigenfunctions from Theorem 5.10.
@@
We distinguish two cases. First, assume $k>1$ and hence $\xi_{k}(\Omega^{\mathrm{ext}})\neq 0$. Let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ be an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})$, extended by zero to $\Omega^{\mathrm{ext}}$ and define
@@
Consider now the case $k=1$. Then $\xi_{1}=0$ and the corresponding eigenfunction is constant. Let $u_{k,R}$ be an eigenfunction associated with $\sigma_{k}^{\mathrm{D}}(\Omega^{\mathrm{ext}}{R})$ normalised to $\|u_{k,R}\|_{L^{2}(\partial\Omega)}=1$. The convergence of the eigenvalue implies
@@
By the Poincaré inequality, this implies that $\left(u_{k,R}\circ\phi\right)-\frac{\int_{\Omega^{*}}u_{k,R}\circ\phi\,\mathrm{d}x}{|\Omega^{*}|}$ converges strongly to zero in $H^{1}(\Omega^{*})$. Moreover, since
$\begin{split}\int_{\partial\Omega^{*}}|u_{k,R}\circ\phi|^{2}\,\mathrm{d}S&\leq\frac{1}{\inf_{x\in\partial\Omega^{*}}|\phi^{\prime}(x)|}\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|u_{k,R}\circ\phi|^{2}\,\mathrm{d}S\\ &=\frac{1}{\inf_{x\in\partial\Omega^{*}}|\phi^{\prime}(x)|}\int_{\partial\Omega}|u_{k,R}|^{2}\,\mathrm{d}S=\frac{1}{\inf_{x\in\partial\Omega^{*}}|\phi^{\prime}(x)|}<\infty,\end{split}$
the average value of $u_{k,R}\circ\phi$ remains bounded. Therefore, there exists a subsequence that converges strongly in $H^{1}(\Omega^{*})$ to a constant $c\in\mathbb{R}$. ∎
Remark 5.12.
An alternative approach to the proof of Theorem 5.10 would be to recognise that ($\mathrm{SN}_{R}$) is isospectral to a mixed Steklov-Neumann problem on $\Omega^{*}\setminus B_{1/R}$ with density $|\phi^{\prime}|$ on $\partial\Omega^{*}$. Since the density does not depend on $R$, and we have the uniform bound (5.4) on the harmonic extension in the small hole, [GirKaLa21, Propositions 4.8 and 4.11] can be used to establish the convergence of the eigenvalues and to prove Corollary 5.11. ∎
@@
The case $k=1$ can be handled as in Corollary 5.11. For $k>1$, we follow the proof of Theorem 5.10. Let $u_{k,R}\in H^{1}\left(\Omega^{\mathrm{ext}}_{R}\right)$ denote an eigenfunction associated to $\sigma_{k}^{\mathrm{N}}\left(\Omega^{\mathrm{ext}}_{R}\right)$ and normalised by $\|\nabla u_{k,R}\|_{L^{2}\left(\Omega^{\mathrm{ext}}_{R}\right)}=1$. Define
@@ 5.2.2 Conformal mapping and Helmholtz equation
In this subsection, we establish the equivalence between the formulation with conformal mappings and the formulation introduced in §3.3, based on the Helmholtz equation. Theorem  5.14 mirrors the convergence result in Theorem 5.9 for the Dirichlet problem.
@@
For the reverse inequality, let $f_{k}\in H^{1}(\Omega^{*})$ be an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$, normalised so that $\int_{\partial\Omega^{*}}|\phi^{\prime}|f_{j}f_{k}\,\mathrm{d}S=\delta_{j,k}$. We construct a cut-off function by
@@
To obtain an upper bound on $\int_{\partial\Omega^{*}}|\phi^{\prime}|\cdot|f_{k,\Lambda}|^{2}\,\mathrm{d}S$, we define the $k$-dimensional spaces
@@ 5.2.3 Conformal mapping and layer potentials
Let $u_{k}\in H^{\frac{1}{2}}(\partial\Omega)$ be an eigenfunction associated with $\tau_{k}(\Omega^{\mathrm{ext}})$, and let $U_{k}$ be its harmonic extension satisfying (1.2). Then, $U_{k}\circ\phi$ coincides, up to a constant factor, with an eigenfunction associated with $\xi_{k}\left(\Omega^{\mathrm{ext}}\right)$.
@@
$\left|f_{k}(x)-w_{k}(0)\right|=\left|w_{k}\left(\frac{x}{|x|^{2}}\right)-w_{k}(0)\right|=\frac{|\nabla w_{k}(x_{0})|}{|x|}$
@@
is a solution of ($\mathrm{CT}$). Note that $W_{k}\in H^{1}(\Omega^{*})$, even though $U_{k}$ is not necessarily in $H^{1}(\Omega^{\mathrm{ext}})$. ∎
@@
Let $n\geq 3$, let $u$ be a solution of ($\mathrm{ES}$) and let $f$ be the restriction of $u$ to $\partial\Omega$. It follows from [AuHa14b, Section 5] that there is a unique $U\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ such that $U=f$ on $\partial\Omega$. Moreover, $U=O(|x|^{2-n})$ by [Xio23, Proposition 12] (see also Proposition 6.1). It follows that UuU-u is harmonic, vanishing on $\Gamma$ and decaying at $\infty$, whence $u=U\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$. So, any solution of ($\mathrm{ES}$) is in $E^{1}\left(\Omega^{\mathrm{ext}}\right)$. It then follows by integration by parts against test function that $u$ is a solution of the finite energy weak Steklov eigenvalue equation ($\mathrm{FE}$). On the other hand, Theorem 3.5 provides a sequence of weak Steklov eigenfunction $u_{k}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ whose restriction to $\partial\Omega$ form an orthogonal basis of $L^{2}(\partial\Omega)$. It follows from [Xio23, Proposition 12] that these functions satisfy (1.2), and hence they are solutions of ($\mathrm{ES}$).
Let $n=2$ and let $u$ be a solution of ($\mathrm{ES}$). Because $u$ is harmonic and bounded, it converges to a finite value at infinity, denoted by $u_{\infty}$. With $\phi$ as in (3.10), we consider the function $v:\Omega^{*}\to\mathbb{R}$, $v=u\circ\phi$ in $\Omega^{*}\setminus\{0\}$ and $v(0)=u_{\infty}$. Then, $v$ is harmonic and satisfies ($\mathrm{CT}$) and so it is a Steklov eigenfunction in the conformal sense. Similarly, any solution of ($\mathrm{CT}$) will give a solution of ($\mathrm{ES}$) once it is inverted through $\phi$. ∎